OpenStudy (anonymous):

Write the following equation in the general form Ax + By + C = 0. 1/2 y - 1/3 x - 1 = 0 -2x + 3y - 6 = 0 2x - 3y + 6 = 0 2x - 3y - 6 = 0

OpenStudy (anonymous):

@matricked

OpenStudy (shubhamsrg):

Multiply both sides by 6.

OpenStudy (anonymous):

Multiply both sides by -6 : -6 * (1/2 y - 1/3 x - 1) = -6 * 0 -3y + 2x + 6 = 0 rearrange x and y terms 2x - 3y + 6 = 0

OpenStudy (shubhamsrg):

Seems like there are 2 options matching.